OpenStudy (anonymous):

Line IJ has an equation of a line y = 3x − 8, and line KL has an equation of a line y = 3x + 4. These two lines are parallel because the slopes of the lines are the same perpendicular because the product of the slopes is −1 both representations of the same line neither parallel nor perpendicular

OpenStudy (anonymous):

@jim_thompson5910 @ganeshie8 @mathmale

jimthompson5910 (jim_thompson5910):

What is the slope of y = 3x − 8? What is the slope of y = 3x + 4?

OpenStudy (anonymous):

idk how to find the slope

jimthompson5910 (jim_thompson5910):

y = mx+b is your general form m is the slope

jimthompson5910 (jim_thompson5910):

example y = 7x+10 y = mx+b so m = 7 is the slope

OpenStudy (anonymous):

ok @jim_thompson5910

jimthompson5910 (jim_thompson5910):

So what is the slope of each?

OpenStudy (anonymous):

7 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

That's the slope of y = 7x+10

jimthompson5910 (jim_thompson5910):

The number in front of the x is the slope but that's NOT the slope of y = 3x - 8

jimthompson5910 (jim_thompson5910):

What's the slope of y = 3x - 8?

OpenStudy (anonymous):

@jim_thompson5910 sorry my computer is running slow and 13

jimthompson5910 (jim_thompson5910):

where are you getting 13?

jimthompson5910 (jim_thompson5910):

are you adding two numbers?

OpenStudy (anonymous):

i multiplied 3 by 7 got 21 and subtracted 8

jimthompson5910 (jim_thompson5910):

the slope is simply the number in front of the x

OpenStudy (anonymous):

@jim_thompson5910 Thank you the answer is a

jimthompson5910 (jim_thompson5910):